. (10pts) Evaluate the integral. So L (6 + 5y?) drdy
Answer #1

si pe (6+593) de dy f $.$(+599) de Joy {! [ (6x + 5y3 x 73 ] ds. ! [ (6y + sy?ty))+(60)4 59 %e= ( 60192 + 5(vs)– (60032 456013 (3+1)-(0 +o)= 4-0=4. tg”(6+93) de des 4